Please Help thanks
Radical 10+x + Radical 11-5x =-1

Answer by lwsshak3(11628)

Simplify:
Radical 10+x + Radical 11-5x =-1
**
√(10+x) + √(11-5x) =-1
√(11-5x)+1=-√(10+x)
square both sides
11-5x+2√(11-5x)+1=10+x
isolate radical
2√(11-5x)=10+x+5x-12=6x-2
square both sides again
4(11-5x)=36x^2-24x+4
44-20x=36x^2-24x+4
36x^2-4x-40